OpenStudy (anonymous):

Rick's automotive chaRges $50 plus $15 for each (15mins) unit of time when making a road call. Twin city repair charges $70 plus $10 for each unit of time. Under what circumstance would it be more economical for a motorist to call Rick's

OpenStudy (anonymous):

let x is a 15 minute unit of time RA = 50 + 15 x TCR = 70 + 10 x we need to know , when is 50 + 15x < 70 + 10 x 5x < 20 x < 4 so when ever the job is less than 4 times 15 minute units (or when the expected time is less than 1 hour)